OpenStudy (anonymous):

Sally is analyzing a circle, y2 + x2 = 9, and a linear function g(x). Will they intersect? x g(x) -2 -1 1 1 4 3

OpenStudy (ranga):

Find the linear function g(x) first. You are given three points on g(x): (-2,-1), (1,1) and (4,3). You just need two points to find g(x). Use the last two points to find g(x).

OpenStudy (anonymous):

These are the answer choices Yes, at positive x-coordinates. Yes, at negative x-coordinates. Yes, at negative and positive x-coordinates. No, they will not intersect. @ranga

OpenStudy (anonymous):

And the screenshot that I had provided shows the two different graphs which I have seen and tried to conclude that the correct answer choice was A.

OpenStudy (ranga):

First locate the points (1,1) and (4,3) on the graph and imagine a line passing through those two points. Where will the line intersect the circle?

OpenStudy (anonymous):

yes @ranga

OpenStudy (ranga):



OpenStudy (ranga):

Look at the x-coordinates of the two points where the line intersects the circle. Are the x-coordinates both positive, both negative, one positive and one negative?

OpenStudy (ranga):



OpenStudy (ranga):

Point A is in the first quadrant where x is positive. Point B is in the third quadrant where x is negative. Therefore, the line intersects the circle at negative and positive x-coordinates.

OpenStudy (anonymous):

Hey do you think you can help me with 2 more problems @ranga

OpenStudy (ranga):

I can try one more.

OpenStudy (anonymous):

How can 1/5x − 2 = 1/3x + 8 be set up as a system of equations? 5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

Ignore the image! @ranga

OpenStudy (ranga):

1/5x − 2 = 1/3x + 8 can be set up as: y = x/5 - 2 and y = x/3 + 8 To get rid of the fraction multiply the first equation throughout by 5 and the second equation throughout by 3: 5y = x - 10 and 3y = x + 24 Bring x to the left hand side: 5y - x = -10 and 3y - x = 24